OpenStudy (anonymous):

Solve the given exponential equation e^2x-3=40

Nnesha (nnesha):

to cancel e you have to take "ln" both side \[\huge\rm e^x \] take ln to cancel e\[\ln e^x =x \] ln and e cancel out so left is just x (example)

Nnesha (nnesha):

\[\huge\rm \color{red}{ln} e^{2x-3} =\color{red}{\ln} 40\] take ln to cancel out e